The company has just hired a new marketing manager who insists that unit sales can be dramatically increased by dropping the selling price from $8 to $7. The marketing manager would like to use the following projections in the budget: 2 Quarter 50,000 65,000 105,000 75,000 85,000 90,000 Budgeted unit sales Selling price per unit $7 per unit 1 Chapter 7: Applying Excel 3 Data 5 Budgeted unit sales Year 2 Quarter Year 3 Quarter 50.000 66,000 10000 75,000 5,000 90,000 7Selling price per unit 8Accounts recevable, beginning balance 9 Sales collected in the quarter sales are made 0 .Sales collected in the quarter 11 Desired ending firished goods inventoryi 12 Finished goods invertiory, beginning 3 Raw materials required to produce one unit 4 Desired ending inventory of raw materials s 5Raw materials inventory, beginning 6 Raw material costs 17Raw materials purchases are paid 18 and 19 Accounts payable for raw materials, beginning balance$81 20 s8 per unit $65,000 75% 25% so% of the budgeted unt sales of the next quarter after sales are made 12,000 units 5 pounds 10% of the next quarters product on needs 23,000 pounds 50.80 per pound 60% Im quarter the purchases are made 40% in the quarter fellowing purchase ,500
media%2Fba5%2Fba52da3f-9d3c-4b5e-bc10-24
Answer #1

A)Quarter 1 = 50000units *7*0.75(current year)= 262500          +      65000 ( opening Accounts Receivable Balance = 327500

Quarter 2 = 350000*0.25 + 65000*0.75*7= 422500

Quarter 3 = 455000*0.25+735000*0.75=113750+551250=655000

Quarter 4 = 75000*0.75*7+735000*0.25= 577500

Total = 1982500 Collection from sale of goods and subsequent and current cash collections

B)

Q1 Q2 Q3 Q4 Q1
Current Quarter sales 50000 65000 105000 75000 85000
Closing inventory 30%
of projected sales of next quarter
19500 31500 22500 25500
less- op inventory -12000 -19500 -31500 -22500
Production per quarter 57500 77000 96000 78000
therefore total equals 308500

c)

Production per quarter 57500 77000 96000 78000 86500
Raw material required 287500 385000 480000 390000 432500
Closing 10% of next quarter 38500 48000 39000 43250
less- opening -23000 -38500 -48000 -39000
Total purchased(pounds) 303000 394500 471000 394250
cost 0.8$ 242400 315600 376800 315400
Total Cost 1250200

D)

Total purchased(pounds) 303000 394500 471000 394250
cost 0.8$ 242400 315600 376800 315400
60% of current quarter 145440 189360 226080 189240
40% of previous Quarter 61500 58176 75744 90432
Total payment 206940 247536 301824 279672
Total Cash Disbursements 1035972

e)Yes since because a set of cash flow assumptions depend on the fact that production would be above 80000units in Q3 and Q4